OpenStudy (sloppycanada):

Is there a difference between solving an equation using the elimination method instead of the substitution method? really bad at the substitution method, much rather use the elimination.

OpenStudy (jack1):

no, no difference in answers, just difference in the approach, s'all

OpenStudy (usukidoll):

elimination method takes out a variable (x or y) and then you can just solve (for x or y) meanwhile substitution means x or y has to be something and then it gets plugged back into the first or second equation... combine like terms solve for x or y lalalalal but both procedures should produce the same results.

OpenStudy (sloppycanada):

Oh okay, so if I solve this question - Solve 4x + 3y = 13 by the substitution method. x + y = 4 using substitution I'll get the same answer?

OpenStudy (usukidoll):

you should get the same answer.

OpenStudy (usukidoll):

trying to think what would be easier.. ah take the second equation and have either x or y by itself

OpenStudy (usukidoll):

take x+y = 4 and rearrange it having either x = ? or y = ?

OpenStudy (jack1):

4x + 4y = 16 4x = 16 - 4y 4x = 13 - 3y 16 - 4y = 13 - 3y is this elimination or am i doing it wrong again?

OpenStudy (usukidoll):

o_O

OpenStudy (usukidoll):

ok I'll give an example where elimination can be used x+y=2 x-y=5 now use elimination method @Jack1

OpenStudy (jack1):

dont give me that look

OpenStudy (jack1):

;P

OpenStudy (jack1):

2-y = 5+y ...? im serious, i dont remember this properly hey

OpenStudy (sloppycanada):

So for this I got x = 1 and y = 3

OpenStudy (jack1):

then solve for y, yeah?

OpenStudy (usukidoll):

for my example x+y=2 x-y=5 add the two equations together @Jack1

OpenStudy (usukidoll):

@sloppycanada that's right when you plug in x =1 and y=3 your equations will be true!

OpenStudy (usukidoll):

4x + 3y = 13 x + y = 4 since x = 1 and y =3 4(1)+3(3) =13 4+9=13 13=13 Similarly 1+3=4 4=4

OpenStudy (jack1):

right, now i remember, cheers @UsukiDoll

OpenStudy (sloppycanada):

Alright so point is - I don't need to use the sub method when it says too.

OpenStudy (sloppycanada):

thanks!